Five workers have been hired to complete a job. If one additional worker is hired, they could complete the job 6 days earlier. If the job needs to be completed 32 days earlier, how many additional workers should be hired?
Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(53996)

.
Five workers have been hired to complete a job.
If one additional worker is hired, they could complete the job 6 days earlier.
If the job needs to be completed 32 days earlier, how many additional workers should be hired?
~~~~~~~~~~~~~~~


                    Solve it step by step.

      Step 1 - determine the number of days needed for 5 workers
               to complete the job.


Let d be the number of days for 5 workers to complete the job.

Then the entire job is 5d worker-days.


6 workers could complete the job in (d-6) days.

Hence, from this perspective, the entire work is  6*(d-6)  worker-days.


It gives us this equation

    5d = 6(d-6),

from which we get

     5d = 6d - 36  --->  36 = 6d - 5d  --->  36 = d.


Hence, 5 workers need 36 days to complete the job,  and the entire job is 5*36 = 180 worker-days.



      Step 2 - determine the number of workers needed 
               to complete the job in 32 days earlier.


The question wants the job be complete in 36-32 = 4 days.


It requires 180/4 = 45 workers.



      Step 3 - determine the number of additional workers 
               to be hired.


The number of additional workers is  45 - 5 = 40.


ANSWER.  40 additional workers should be hired to complete the job in 32 days earlier.

Solved.



Answer by math_tutor2020(3844)


Answer: 40 extra workers

Explanation

Let's say the job is to move 9000 boxes.
I'm picking this value since it's a multiple of 5 and 6.
You can change 9000 to anything else you want to get the same answer at the end.

If there are 5 workers, then each worker handles 9000/5 = 1800 boxes.
Each worker has a unit rate of 1800/x boxes per day.
x is the number of days to finish the job with 5 workers.
rate = amountDone/time

When there are 6 workers, each person now gets 9000/6 = 1500 boxes.
The unit rate per person is 1500/(x-6) where the x-6 refers to finishing the job 6 days early compared to the previous scenario.

If we assume each worker has the same unit rate, then we get this equation
1800/x = 1500/(x-6)
Solving it leads to x = 36.
I'll let the student handle the scratch work.

It takes 36 days for 5 workers to do the job.
It takes 30 days for 6 workers to do the same job.
Each worker moves 50 boxes per day because 1800/36 = 50 or 1500/30 = 50.

n = number of additional workers to hire
n+5 = total number of workers when including the original 5 workers
9000/(n+5) = number of boxes each worker handles
The goal is to finish 32 days early. So instead of taking 36 days it should take 36-32 = 4 days

rate*time = amount done
(50 boxes per day)*(4 days) = 9000/(n+5) boxes
50*4 = 9000/(n+5)
Solving for n leads to n = 40
I'll let the student handle the scratch work.

More practice with a similar question is found here